# Relaxed Equivariance via Multitask Learning

### v1

@@ 1 Introduction
Equivariant machine learning models have achieved notable success across various domains, such as computer vision (Weiler et al., 2018; Yu et al., 2022), dynamical systems (Han et al., 2022; Xu et al., 2024), chemistry (Satorras et al., 2022; Brandstetter et al., 2022), and structural biology (Jumper et al., 2021). For example, incorporating equivariance w.r.t. translations and rotations ensures the correct handling of complex structures like graphs and molecules (Schütt et al., 2021; Bronstein et al., 2021; Thölke & Fabritiis, 2022; Liao et al., 2024). Equivariant machine learning models benefit from this inductive bias by explicitly leveraging symmetries of the data during the architecture design. Typically, such architectures have highly constrained layers with restrictions on the form and action of weight matrices and nonlinear activations (Batzner et al., 2022; Batatia et al., 2023). This may come at the expense of higher computational cost, making it sometimes challenging to scale equivariant architectures, particularly those relying on spherical harmonics and irreducible representations (Thomas et al., 2018; Fuchs et al., 2020; Liao & Smidt, 2023; Luo et al., 2024).
Equivariant layers are not the only way to incorporate symmetries into deep neural networks. Several approaches have been proposed to either offload the equivariance restrictions to faster networks (Kaba et al., 2022; Baker et al., 2024; Ma et al., 2024) or simplify the constraints by introducing averaging operations (Puny et al., 2022; Duval et al., 2023; Lin et al., 2024; Huang et al., 2024). Nonetheless, these directions cannot seamlessly leverage unconstrained architectures that do not bake symmetries into their design by simply altering the training protocol. To this aim, a widely adopted strategy to replace ‘hard’ equivariance (i.e., built into the architecture itself) with a ‘soft’ one, is data augmentation (Quiroga et al., 2019; Bai et al., 2021; Gerken et al., 2022; Iglesias et al., 2023; Xu et al., 2023; Yang et al., 2024), whereby the training protocol of an arbitrary (unconstrained) network is augmented by assigning the same label to group orbits (e.g., rotated and translated versions of the input). In fact, recent works have shown that unconstrained architectures may offer a valid alternative provided that enough data are available (Wang et al., 2024; Abramson et al., 2024).
@@
In this work, we present REMUL: Relaxed Equivariance via Multitask Learning. REMUL is a training procedure that aims to learn approximate equivariance during training for unconstrained networks using a multitask approach with adaptive weights. We conduct a comprehensive evaluation of unconstrained models trained with REMUL, comparing their performance and computational efficiency against equivariant models. We consider Transformers and Graph Neural Networks (GNNs), as well as their equivariant versions, as our main baselines.
@@ 4 Quantifying Learned Equivariance
Using group transformations to measure and assess the symmetries of ML models has been studied in several domains (Lyle et al., 2020; Kvinge et al., 2022; Moskalev et al., 2023; Gruver et al., 2024; Speicher et al., 2024). Inspired by the idea of frame-averaging (Puny et al., 2022; Duval et al., 2023; Lin et al., 2024), in this section, we introduce a metric to quantify the degree of equivariance exhibited by a function $f$.
Starting from Equation 1, and assume that $G$ is a finite or compact group with normalized Haar measure $\mu$, the group integration of both sides yields:
@@
When $G$ is a large or continuous group, we approximate the integrals using a Monte Carlo approach with samples $\{g_{i}\}_{i=1}^{M}$ from $G$:
@@
$E(f,G)=\frac{1}{|D|}\sum_{x\in D}\left\|\frac{1}{M}\sum_{i=1}^{M}\rho(g_{i})(f(x))-\frac{1}{M}\sum_{i=1}^{M}f(\phi(g_{i})(x))\right\|$ (9)
This error indicates the average deviation of a function $f$ from perfect equivariance across the data distribution $D$ (lower value means more equivariant function).
@@
$E^{\prime}(f,G)=\frac{1}{|D|}\sum_{x\in D}\frac{1}{M}\sum_{i=1}^{M}\left\|f(\phi(g_{i})(x))-\rho(g_{i})(f(x))\right\|$ (10)
Equation 9 & Equation 10 indicate a practical metric for evaluating how closely the function $f$ approximates perfect equivariance throughout a data distribution $D$ (which should be zero for a perfect equivariance function). We noticed that both measures have very similar behavior in our experiments, where $E$ and $E^{\prime}$ are near zero for equivariant models. We also demonstrate that increasing the value of $\beta$ in Equation 7 results in a less equivariant error for $E$ and $E^{\prime}$.
@@ 5 Related Work
Equivariant ML Models. In the vision domain, group convolutions have proven to be a powerful tool for handling rotation equivariance for images and enhanced model generalization (Cohen & Welling, 2016; Cohen et al., 2019; Weiler & Cesa, 2021; Qiao et al., 2023). Similarly, the development of equivariant architectures with respect to roto-translations for geometric data has been an active area of research (Chen et al., 2021a; Satorras et al., 2022; Han et al., 2022; Xu et al., 2024). Techniques that use spherical harmonics and irreducible representations have shown a large success in modeling graph-structured data, such as SE(3)-Transformers (Fuchs et al., 2020), Tensor Field Networks (Thomas et al., 2018), and DimeNet (Gasteiger et al., 2022). More recently, Brehmer et al. (2023) introduced an E(3) equivariant Transformer that employs geometric algebra for processing 3D point clouds.
@@
Data Augmentation and Unconstrained Models. Alternatively, integrating transformations through data augmentation is a widely used strategy across multiple vision tasks, enhancing performance in image classification (Perez & Wang, 2017; Inoue, 2018; Rahat et al., 2024), object detection (Zoph et al., 2019; Wang et al., 2019; Kisantal et al., 2019), and segmentation (Negassi et al., 2021; Chen et al., 2021b; Yu et al., 2024). For geometric data, Hu et al. (2021) has adapted a Graph Neural Network architecture with data augmentation to process 3D molecular structures. In parallel, Dosovitskiy et al. (2021) introduced that Vision Transformers (ViTs) with a large amount of training data can achieve comparable performance with Convolutional Neural Networks (CNNs), obviating the need for explicit translation equivariance within the architecture. Recently, this has shown to be effective for handling geometric data (Wang et al., 2024; Abramson et al., 2024).
Learning Symmetries and Approximate Equivariance. Several studies have shown that the layers of CNN architectures can be approximated for a soft constraint (Wang et al., 2022; van der Ouderaa et al., 2022; Veefkind & Cesa, 2024; Wu et al., 2024). Conversely, van der Ouderaa et al. (2023) extends the Bayesian model selection approach to learning symmetries in image datasets. For geometric data, recent works have modified the EGNN framework (Satorras et al., 2022) using an invariant function (Zheng et al., 2024) or adversarial training procedure (Yang et al., 2023).
@@ 6 Experiments and Discussion
We evaluate our method on different tasks for geometric data: N-body dynamical system (Section 6.1), motion capture (Section 6.2), and molecular dynamics (Section 6.3). For unconstrained models, we apply REMUL to Transformers and Graph Neural Networks. We then compare against their equivariant baselines: SE(3)-Transformer (Fuchs et al., 2020), Geometric Algebra Transformer (Brehmer et al., 2023), and Equivariant Graph Neural Networks (Satorras et al., 2022) as well as unconstrained models with data augmentation. We consider learning the rotation group $SO(3)$ for REMUL and data augmentation and we subtract the center of mass for translation. We also conduct a comparative analysis for the computational requirements of unconstrained models and equivariant models in Section 6.4. Lastly, we discuss the loss surfaces in Section 6.5. Implementation details and additional experiments can be found in Appendix A & Appendix B.
@@ 6.2 Motion Capture
Table 1: Performance on Motion Capture dataset: MSE (×$\times 10^{-2}$). REMUL procedure and data augmentation were applied to standard Transformer. First, Second (highlighted). REMUL comes the best in both tasks. SE(3)-Transformer and GATr go worse than Transformer.
@@
We apply our training procedure REMUL to the Transformer architecture and compare it with SE(3)-Transformer, Geometric Algebra Transformer (GATr), standard Transformer, and Transformer trained with data augmentation. Our results are presented in Table 1. For REMUL, we also provide plots on how the performance and equivariance error change w.r.t. the penalty parameter $\beta$ in Figure 3.
Table 1 indicates that when processing 3D positions related to human motions, both SE(3)-Transformer and GATr perform worse than the standard Transformer. This outcome is noteworthy because human motion inherently lacks symmetry along the vertical or gravity axis. Consequently, the assumption of equivariance across all axes may not be beneficial or even detrimental. In contrast, a standard Transformer trained with REMUL has the best performance in both tasks. Following Figure 3, there is a noticeable trade-off in model performance with different values of penalty parameter $\beta$. Best performance is observed at an intermediate level of equivariance, where the model balances between being too rigid (fully equivariant) and too flexible (non-equivariant). This finding underscores the importance of carefully considering the specific characteristics of the data and the task when designing equivariant architectures.
6.3 Molecular Dynamics
@@
From the results presented in Table 2, GNN trained with REMUL outperforms EGNN in six out of eight molecules. Interestingly, a standard GNN, without data augmentation or REMUL, surpasses the performance of EGNN for two molecules: Aspirin and Toluene. In Figure 4 & Figure 8, we observe that the optimal performance of each molecule is attained at different values of the penalty parameter $\beta$. For instance, Malonaldehyde exhibits a direct correlation between model performance and equivariance, where a higher $\beta$ yields better performance. Conversely, for most other molecules, there appears to be a pronounced trade-off where the best performance is achieved at a lower value of $\beta$. This is particularly evident with molecules like Aspirin, where a standard GNN architecture outperforms EGNN.
We also plot the 3D structures of the eight molecules in Figure 10. Molecules such as Malonaldehyde, characterized by their symmetric components, might be ideally suited for equivariant design. However, this advantage does not apply to all molecules. Aspirin on the other side, might have an asymmetric structure and exhibit a range of interactions and dynamic states that equivariant models might simplify. Consequently, for such molecules, less equivariant models could potentially offer more accurate predictions.
@@ 6.5 Loss Surface
We visualize the loss surface of the Geometric Algebra Transformer (GATr) and Transformer in Figure 1, both trained on the N-body dynamical system. We observe that the Transformer architecture exhibits a more favorable loss shape around its local minima, characterized by a convex structure. This might suggest that the optimization path for the Transformer is smoother and potentially easier to navigate during training, leading to more stable convergence. In contrast, the loss surface of GATr appears more erratic and rugged. This complexity in the loss landscape can indicate multiple local minima and a higher sensitivity to initial conditions or parameter settings. Such characteristics might complicate the training process, requiring more careful tuning of hyperparameters. We leave this for future work to analyze in more detail how the optimization path for each model behaves during training.
@@ A.2 Motion Capture
Motion Capture dataset by CMU (2003) features 3D trajectory data that records a range of human motions, and the task involves predicting the final trajectory based on initial positions and velocities. We have reported results for two types of motion: Walking (Subject #$35$) and Running (Subject #$9$). Following the standard experimental setup in the literature on this task (Han et al., 2022; Huang et al., 2022; Xu et al., 2024), we apply a train/validation/test split of $200$/$600$/$600$ for Walking and $200$/$240$/$240$ for Running. The interval between trajectories, $\Delta T=30$ for both tasks. For model hyperparameters, we fine-tuned around the same in Table 3 and found it the best for each model except for the Geometric Algebra Transformer we increased the attention blocks to $12$. We train each model for $2000$ epochs with batch size $=12$ (outlined in Table 4). For REMUL, $\alpha=1$.

### v2

@@ 1 Introduction
Equivariant machine learning models have achieved notable success across various domains, such as computer vision (Weiler et al., 2018; Yu et al., 2022), dynamical systems (Han et al., 2022; Xu et al., 2024), chemistry (Satorras et al., 2021; Brandstetter et al., 2022), and structural biology (Jumper et al., 2021). For example, incorporating equivariance w.r.t. translations and rotations ensures the correct handling of complex structures like graphs and molecules (Schütt et al., 2021; Bronstein et al., 2021; Thölke & Fabritiis, 2022; Liao et al., 2024). Equivariant machine learning models benefit from this inductive bias by explicitly leveraging symmetries of the data during the architecture design. Typically, such architectures have highly constrained layers with restrictions on the form and action of weight matrices and nonlinear activations (Batzner et al., 2022; Batatia et al., 2022). This may come at the expense of higher computational cost, making it sometimes challenging to scale equivariant architectures, particularly those relying on spherical harmonics and irreducible representations (Thomas et al., 2018; Fuchs et al., 2020; Liao & Smidt, 2023; Luo et al., 2024). On the other hand, equivariance constraints might limit the expressive power of the network, restricting its ability to act as a universal architecture (Dym & Maron, 2021; Joshi et al., 2023).
Equivariant layers are not the only way to incorporate symmetries into deep neural networks. Several approaches have been proposed to either offload the equivariance restrictions to faster networks (Kaba et al., 2022; Mondal et al., 2023; Baker et al., 2024; Ma et al., 2024; Panigrahi & Mondal, 2024) or simplify the constraints by introducing averaging operations (Puny et al., 2022; Duval et al., 2023; Lin et al., 2024; Huang et al., 2024). Nonetheless, while these approaches leverage unconstrained architectures, they often require additional networks or averaging techniques to achieve equivariance and may not rely solely on adjustments to the training protocol. To this aim, a widely adopted strategy to replace ‘hard’ equivariance (i.e., built into the architecture itself) with a ‘soft’ one, is data augmentation (Quiroga et al., 2019; Bai et al., 2021; Gerken et al., 2022; Iglesias et al., 2023; Xu et al., 2023; Yang et al., 2024), whereby the training protocol of an arbitrary (unconstrained) network is augmented by assigning the same label to group orbits (e.g., rotated and translated versions of the input). In fact, recent works have shown that unconstrained architectures may offer a valid alternative provided that enough data are available (Wang et al., 2024; Abramson et al., 2024).
@@
In this work, we present REMUL: Relaxed Equivariance via Multitask Learning. REMUL is a training procedure that aims to learn approximate equivariance during training for unconstrained networks using a multitask approach with adaptive weights. We conduct a comprehensive evaluation of unconstrained models trained with REMUL, comparing their performance and computational efficiency against equivariant models. We consider Transformers and Graph Neural Networks (GNNs), as well as their E(3) equivariant versions, as our main baselines, focusing on roto-translation group.
@@ 4 Quantifying Learned Equivariance
Using group transformations to measure and assess the symmetries of ML models has been studied in several domains (Lyle et al., 2020; Kvinge et al., 2022; Moskalev et al., 2023; Gruver et al., 2023; Speicher et al., 2024). Inspired by the idea of frame-averaging (Puny et al., 2022; Duval et al., 2023; Lin et al., 2024), in this section, we introduce a metric to quantify the degree of equivariance exhibited by a function $f$.
Starting from Equation 1, the group integration of both sides w.r.t. the normalized Haar measure $\mu$ yields:
@@
When $G$ is a large or continuous group, as is the case in our work, the integrals over $G$ may not be computable in closed form. Therefore, we approximate the integrals using a Monte Carlo approach with samples $\{g_{i}\}_{i=1}^{M}$ from $G$:
@@
$E(f,G)=\frac{1}{|D|}\sum_{x\in D}\left\|\frac{1}{M}\sum_{i=1}^{M}\rho(g_{i})(f(x))-\frac{1}{M}\sum_{i=1}^{M}f(\phi(g_{i})(x))\right\|_{2}$ (12)
Here $\|\cdot\|_{2}$ denotes an $L_{2}$ norm (for non-scalar function). This error indicates the average deviation of a function $f$ from perfect equivariance across the data distribution $D$ (lower value means more equivariant function).
@@
$E^{\prime}(f,G)=\frac{1}{|D|}\sum_{x\in D}\frac{1}{M}\sum_{i=1}^{M}\left\|f(\phi(g_{i})(x))-\rho(g_{i})(f(x))\right\|_{2}$ (13)
Equation 12 & Equation 13 indicate a practical metric for evaluating how closely the function $f$ approximates perfect equivariance throughout a data distribution $D$ (which should be zero for a perfect equivariance function). In practice, we use $M=100$ samples from the group and noticed this was sufficient to obtain stable results. We also observed that both measures have very similar behavior in our experiments, where $E$ and $E^{\prime}$ are near zero for equivariant models. We also demonstrate that increasing the value of $\beta$ in Equation 7 results in a less equivariant error for $E$ and $E^{\prime}$.
@@ 5 Related Work
Equivariant ML Models. In the vision domain, group convolutions have proven to be a powerful tool for handling rotation equivariance for images and enhanced model generalization (Cohen & Welling, 2016; Cohen et al., 2019; Weiler & Cesa, 2019; Qiao et al., 2023). Similarly, the development of equivariant architectures with respect to roto-translations for geometric data has been an active area of research (Chen et al., 2021a; Satorras et al., 2021; Han et al., 2022; Xu et al., 2024). Techniques that use spherical harmonics and irreducible representations have shown a large success in modeling graph-structured data, such as SE(3)-Transformers (Fuchs et al., 2020), Tensor Field Networks (Thomas et al., 2018), and DimeNet (Gasteiger et al., 2020). More recently, Brehmer et al. (2023) introduced an E(3) equivariant Transformer that employs geometric algebra for processing 3D point clouds.
@@
Data Augmentation and Unconstrained Models. Alternatively, integrating transformations through data augmentation is a widely used strategy across multiple vision tasks, enhancing performance in image classification (Perez & Wang, 2017; Inoue, 2018; Rahat et al., 2024), object detection (Zoph et al., 2020; Wang et al., 2019; Kisantal et al., 2019), and segmentation (Negassi et al., 2022; Chen et al., 2021b; Yu et al., 2023). For geometric data, Hu et al. (2021) has adapted a Graph Neural Network architecture with data augmentation to process 3D molecular structures. In parallel, Dosovitskiy et al. (2021) introduced that Vision Transformers (ViTs) with a large amount of training data can achieve comparable performance with Convolutional Neural Networks (CNNs), obviating the need for explicit translation equivariance within the architecture. Recently, this has shown to be effective for handling geometric data (Wang et al., 2024; Abramson et al., 2024).
Learning Symmetries and Approximate Equivariance. Several studies have shown that the layers of CNN architectures can be approximated for a soft constraint (Wang et al., 2022; van der Ouderaa et al., 2022; Romero & Lohit, 2022; Veefkind & Cesa, 2024; Wu et al., 2024; McNeela, 2023). Conversely, van der Ouderaa et al. (2023) extends the Bayesian model selection approach to learning symmetries in image datasets. Yeh et al. (2022) introduced a parameter-sharing scheme to achieve permutations and shifts equivariances in Gaussian distributions. Recent works have relaxed the hard constrained models to a soft constraint by adding unconstrained layers in the architecture design (Finzi et al., 2021a; Pertigkiozoglou et al., 2024), canonicalization network (Lawrence et al., 2024). , or explicit relaxation (Kaba & Ravanbakhsh, 2023). Additionally, Lin et al. (2019) modified the loss of CNN for segmentation task. Shakerinava et al. (2022) introduced a method to learn equivariant representation using the group invariants, while Bhardwaj et al. (2023) defined a regularizer that injects the equivariance in the latent space of the network by explicitly modeling transformations with additional learnable maps. In contrast, several works have started from pre-trained models (Basu et al., 2023; Kim et al., 2023b). Furthermore, the EGNN framework (Satorras et al., 2021) has been modified using an invariant function (Zheng et al., 2024) or adversarial training procedure (Yang et al., 2023). However, in our work, we start from completely unconstrained models, without imposing any equivariance constraints on the space of functions within the architecture. Moreover, we didn’t assume a specific class of models or introduce additional parameters, which increases the applicability of our method to various domains and makes it computationally efficient.
@@ 6 Experiments and Discussion
We evaluate our method on different tasks for geometric data: N-body dynamical system (Section 6.1), motion capture (Section 6.2), and molecular dynamics (Section 6.3). For unconstrained models, we apply REMUL to Transformers and Graph Neural Networks. We then compare against their equivariant baselines: SE(3)-Transformer (Fuchs et al., 2020), Geometric Algebra Transformer (Brehmer et al., 2023), and Equivariant Graph Neural Networks (Satorras et al., 2021) as well as unconstrained models with data augmentation. We consider learning the rotation group $SO(3)$ for REMUL and data augmentation and we subtract the center of mass for translation. We use the equivariance metric defined in Equation 12 to analyze our results. We also conduct a comparative analysis for the computational requirements of unconstrained models and equivariant models in Section 6.4. Lastly, we discuss the loss surfaces in Section 6.5. Implementation details and additional experiments can be found in Appendix A & Appendix B.
@@ 6.1 N-Body Dynamical System
Table 1: Performance on Motion Capture dataset: MSE (×$\times 10^{-2}$). REMUL procedure and data augmentation were applied to standard Transformer & MLP. First, Second (highlighted). REMUL comes the best in both tasks.
@@
EMLP RPP PER MLP Data Augmentation REMUL
Walking (Subject #$35$) $7.01$±$\pm 0.46$ $6.99$±$\pm 0.21$ $7.48$±$\pm 0.39$ $6.80$±$\pm 0.18$ $\bf 6.37$±$\pm 0.04$ $\bf 6.04$±$\pm 0.09$
Running (Subject #$9$) $57.38$±$\pm 8.39$ $34.18$±$\pm 2.00$ $\bf 33.03$±$\pm 0.37$ $39.56$±$\pm 2.25$ $40.23$±$\pm 0.94$ $\bf 32.57$±$\pm 1.47$
6.2 Motion Capture
@@
We apply our training procedure REMUL to the Transformer architecture and compare it with SE(3)-Transformer, Geometric Algebra Transformer (GATr), standard Transformer, and Transformer trained with data augmentation. We also compare with Equivariant MLP (Finzi et al., 2021b), as well as two approximate equivariance architectures: Residual Pathway Priors (RPP) (Finzi et al., 2021a), and Projection-Based Equivariance Regularizer (PER) (Kim et al., 2023a). As these architectures are designed specifically on MLP and linear layers, we apply our method to a standard MLP with a similar number of parameters. Our results are presented in Table 1. For REMUL, we provide plots on how the performance and equivariance error change w.r.t. the penalty parameter $\beta$ in Figure 3.
@@
Table 1 indicates that when processing 3D positions related to human motions, both SE(3)-Transformer and GATr perform worse than the standard Transformer. This outcome is noteworthy because human motion inherently lacks symmetry along the vertical or gravity axis. Consequently, the assumption of equivariance across all axes may not be beneficial or even detrimental. In contrast, a standard Transformer trained with REMUL has the best performance in both tasks.
Following Figure 3, there is a noticeable trade-off in model performance with different values of penalty parameter $\beta$. Best performance is observed at an intermediate level of equivariance, where the model balances between being too rigid (fully equivariant) and too flexible (non-equivariant). This finding underscores the importance of carefully considering the specific characteristics of the data and the task when designing equivariant architectures.
6.3 Molecular Dynamics
From the results presented in Table 2, GNN trained with REMUL outperforms EGNN in six out of eight molecules. Interestingly, a standard GNN, without data augmentation or REMUL, surpasses the performance of EGNN for two molecules: Aspirin and Toluene. In Figure 4 & Figure 9, we observe that the optimal performance of each molecule is attained at different values of the penalty parameter $\beta$. For instance, Malonaldehyde exhibits a direct correlation between model performance and equivariance, where a higher $\beta$ yields better performance. Conversely, for most other molecules, there appears to be a pronounced trade-off where the best performance is achieved at a lower value of $\beta$. This is particularly evident with molecules like Aspirin, where a standard GNN architecture outperforms EGNN. We also plot the 3D structures of the eight molecules in Figure 11. Molecules such as Malonaldehyde, characterized by their symmetric components, might be ideally suited for equivariant design. However, this advantage does not apply to all molecules. Aspirin on the other side, might have an asymmetric structure and exhibit a range of interactions and dynamic states that equivariant models might simplify. Consequently, for such molecules, less equivariant models could potentially offer more accurate predictions.
@@ 6.5 Loss Surface
We visualize the loss surface of the Geometric Algebra Transformer (GATr) and Transformer in Figure 1, trained on the N-body dynamical system. We observe that the Transformer architecture exhibits a more favorable loss shape around its local minima, characterized by a convex structure. This might suggest that the optimization path for the Transformer is smoother and potentially easier to navigate during training, leading to more stable convergence. In contrast, the loss surface of GATr appears more erratic and rugged. This complexity in the loss landscape can indicate multiple local minima and a higher sensitivity to initial conditions or parameter settings. Such characteristics might complicate the training process, requiring more careful tuning of hyperparameters. We leave this for future work to analyze how the optimization path for each model behaves during training.
@@ A.2 Motion Capture
Motion Capture dataset by CMU (2003) features 3D trajectory data that records a range of human motions, and the task involves predicting the final trajectory based on initial positions and velocities. We have reported results for two types of motion: Walking (Subject #$35$) and Running (Subject #$9$).
@@
Hyperparameters Equivariant MLP RPP PER standard MLP
#hidden dim $532$ $348$ $532$ $680$
#layers $3$ $3$ $3$ $3$
Following the standard experimental setup in the literature on this task (Han et al., 2022; Huang et al., 2022; Xu et al., 2024), we apply a train/validation/test split of $200$/$600$/$600$ for Walking and $200$/$240$/$240$ for Running. The interval between trajectories, $\Delta T=30$ for both tasks. For model hyperparameters, we fine-tuned around the same in Table 3 and found it the best for each model except for the Geometric Algebra Transformer we increased the attention blocks to $12$. We train each model for $2000$ epochs with batch size $=12$. For MLP comparisons, all models and baselines have the same number of layers and parameters. More details in Table 4. For REMUL, $\alpha=1$.